OpenStudy (anonymous):

URGENT! To open these doors, you must match the number and type of solutions for the following two functions in standard form. f(x) = x2 + 6x – 16 g(x) = x2 +6x + 1 Match the following descriptions of the solutions to each of the functions above. Hint: they each have their own match. Two real irrationals solutions Two real rationals solutions

OpenStudy (anonymous):

x^2

OpenStudy (one098):

\[x^{2}\]

OpenStudy (anonymous):

yes, that! thank you

OpenStudy (one098):

@Joel_the_boss

OpenStudy (cwrw238):

find the value of the discriminant b^2 - 4ac the a,b and c refer to the constants in the standard form ax^2 + bx + c = 0

OpenStudy (anonymous):

I honestly have no idea what any of that means. I'm on the verge of flunking math so.. do I just substitute in random numbers?

OpenStudy (cwrw238):

compare the frirst of your equations to the above atandard form a x^2 + bx + c = 0 x^2 + 6x - 16 = 0 a = 1, b = 6 and c = -16 do you follow?

OpenStudy (anonymous):

rogger that

OpenStudy (cwrw238):

right so now you plug these values into b^2 - 4ac b^2 - 4ac = 6^2 - 4 * 1 * -16 = 36 + 64 = 100

OpenStudy (anonymous):

okayyy

OpenStudy (cwrw238):

so the value is 100 which is a perfect square which indicates that the equation has two real rational solutions

OpenStudy (anonymous):

gotchya

OpenStudy (cwrw238):

now if you do the same with the second equation you ought to find that the value is a positive number but not a perfect sqaure meaning that there are 2 real irrational solutions

OpenStudy (anonymous):

Wow, you are so lovely! thank you a million zillion times! you saved my life and you don't even know it<3 :)

OpenStudy (cwrw238):

yw

OpenStudy (cwrw238):

the other 2 values of the discriminant negative - means the roots are non real (complex) zero - roots are real and duplicate

OpenStudy (anonymous):

Got it, thanks again :)